# Explainability of Machine Learning Models under Missing Data

### v1

@@ 5.2 Global feature importance plot
Regardless of the missing rates, the global feature importance consistently highlights three main predictors of housing prices: Latitude, Longitude, and Median Income (MedInc). However, their orders seem to change compared to the plot on the original data. For example, at a missing rate of 0.2, the plot on the original data highlights Lattitude the most, then Longitude, and then MedInc. However, all the remaining methods highlight MedInc the most.
@@
Interestingly, across all missing rates, Xgboost produces the plots that deviate the most from the ground truth for the MedInc, Lattitude, and Longitude features. Specifically, for the XGBoost model, besides presenting high mean absolute Shapley values for these three features with a range of $(+0.28)$ to $(+0.66)$, an interesting difference is observed where MedInc is recorded as more important than Latitude and Longitude, underscoring its critical role in predicting housing prices. Meanwhile, for the five imputation methods, those data have a similar distribution to the results on the original data but at lower, with a range of $[+0.05;+0.22]$.
Additionally, the figures also show that when the missing rate increases from 20% to 40%, the Shapley values for these three comparison features generally decline, while XGBoost experiences an upward trend. The opposite pattern is seen when the missing rate rises to 60%. At the highest missing rate, there is a slight variation in these data, with an amplitude of about ±$\pm 0.03$.
(a) Original
(b) Xgboost
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(a) Original
(b) Xgboost
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(a) Original
(b) Xgboost
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(a) Original
(b) Xgboost
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
5.3 Beeswarm plot analysis
5.3.1 Beeswarm plot for the California datasets
The beeswarm plot for the California datasets at missing rates $0.2,0.4,0.6,$ and $0.8$ are presented in Figures 5, 6, 7, 8, respectively. In these figures, we focus on three key features (Latitude, Longitude, and Median Income - MedInc) because the others show little impact.
XGBoost seems to be stable, MedInc consistently shows the highest Shapley values, ranging from (0.5(-0.5( - 0.5) to around (+1.5(+1.5( + 1.5) across all missing rates, emphasizing its strong influence on the model outputs. The imputation methods show similar patterns with Shapley values for key features slightly reduced compared to the XGBoost model, but they still highlight the importance of Latitude, Longitude, and MedInc in the predictive modeling.
Considering the results at each missing rate, we observed similar distributions in the Shapley values for Latitude, Longitude, and MedInc on both the original and imputed data using various imputation methods, generally fluctuating around zero. However, XGBoost shows a distribution that is more skewed from zero. For example, at a missing rate of 20% in figure 5, the Shapley values in the original data for Latitude and Longitude range from $(-0.75)$ to $(+0.75)$, and for MedInc from $(-0.5)$ to $(+0.7)$. Likewise, the ranges for models with imputation methods are $[-0.6;0.6]$ and $[-0.4;0.6]$, respectively, indicating that the imputation does not significantly alter the feature importance. Meanwhile, XGBoost displays a different distribution from $(-1.0)$ to $(+0.5)$ for Latitude and Longitude and a wider range from $(-0.5)$ to $(+1.5)$ for MedInc.
(a) Original
@@ 5.3.1 Beeswarm plot for the California datasets
5.3.2 Beeswarm plot for the diabetes datasets
The beeswarm plot for the diabetes datasets at missing rates $0.2,0.4,0.6,$ and $0.8$ are presented in Figures 9, 6, 11, 12, respectively. In these figures, the trend regarding the width of the Shapley values range is quite similar across the models, while reverse patterns can be observed in their distribution. When the missing rate increases, the range of Shapley values in the XGBoost model narrows down even though it starts with a width of approximately 0.9. Similarly, the range in the other models shows a slight reduction from about [-0.4; 0.4] at a 20% missing rate to [-0.2; 0.2] at the highest missing rate.
Regarding the distribution of Shapley values, across all missing rates, the original and imputed data with imputation methods exhibit a relatively uniform distribution around the zero value. In contrast, the XGBoost model shows an increasingly skewed distribution as the missing rate rises. Notably, when the missing rate increases from 40% onwards, the Shapley values for features such as “bmi,” “bp,” “s5,” “age,” and “s6” tend to be positive, while the values for other features are negative. Another difference is that the SHAP value for the feature “s4” mostly shows positive values but turns negative when the missing rate reaches 80%.
(a) Original
@@ 5.3.2 Beeswarm plot for the diabetes datasets
5.4 MSE analysis
$r$ criteria Xgboost MI MICE DIMV missForest SOFT
0.2 MSE 0.116 0.136 0.124 0.124 0.089 0.139
MSE Shap 0.184 0.092 0.035 0.028 0.037 0.049
0.4 MSE 0.447 0.473 0.514 0.503 0.445 0.497
MSE Shap 0.746 0.244 0.166 0.139 0.177 0.196
0.6 MSE 0.114 0.133 0.122 0.125 0.089 0.137
MSE Shap 0.183 0.098 0.033 0.026 0.044 0.049
0.8 MSE 0.115 0.132 0.124 0.123 0.090 0.139
MSE Shap 0.160 0.091 0.039 0.030 0.031 0.054
Table 2: MSE on the test set and MSE of Shapley values on the California dataset. Here, MI denotes Mean Imputation, and SOFT denotes SOFT-IMPUTE. The boil indicates the best performance.
The performance of various imputation methods was evaluated for the California dataset with different data missing rates in Table 2, and for the diabetes dataset in Table 3. Recall that here, the MSE is the MSE of the true label and predicted values on the test set. Also, MSE Shap is the MSE between the Shapley values of predicted and true labels on the test set as well.
From Table 2, one can see that across all missing data rates, missForest consistently achieved the lowest MSE values, indicating superior performance in terms of imputation accuracy. Specifically, for $r=0.2$, missForest achieved an MSE of 0.089, significantly lower than the second-best method, Xgboost, with an MSE of 0.116. For higher missing data rates ($r=0.4,0.6,0.8$), missForest maintained its leading position with MSE values of 0.445, 0.089, and 0.090, respectively, outperforming other methods such as MI, MICE, and DIMV.
Also in Table 2, we can see that in terms of MSE Shapley values, DIMV demonstrated the best performance, indicating its robustness in maintaining the interpretability of the imputed data. For $r=0.2$, DIMV achieved an MSE Shap of 0.028, outperforming MICE (0.035) and missForest (0.037). This trend continued across all missing data rates, with DIMV recording the lowest MSE Shap values of 0.139, 0.026, and 0.030 for $r=0.4,0.6,$ and $0.8$, respectively. We, therefore, observe that there are different imputation methods that perform the best in terms of prediction performance and in terms of explainability, underscoring the importance of selecting appropriate imputation methods based on specific criteria in the context of missing data analysis.
$r$ criteria Xgboost MI MICE DIMV missForest SOFT
0.2 MSE 0.209 0.199 0.169 0.177 0.174 0.187
MSE Shap 0.058 0.020 0.015 0.016 0.018 0.017
0.4 MSE 0.293 0.284 0.307 0.280 0.298 0.285
MSE Shap 0.127 0.028 0.029 0.026 0.030 0.026
0.6 MSE 0.407 0.399 0.423 0.399 0.480 0.407
MSE Shap 0.156 0.038 0.041 0.038 0.042 0.039
0.8 MSE 0.527 0.509 0.569 0.549 0.683 0.562
MSE Shap 0.215 0.047 0.051 0.052 0.059 0.050
Table 3: MSE on the test set and MSE of Shapley values on the diabetes dataset. Here, MI denotes Mean Imputation, and SOFT denotes SOFT-IMPUTE. The boil indicates the best performance.
Focusing on the results with the diabetes dataset in Table 3, overall, both MSE and MSE Shap experience a significant increase when the missing rate increases. In particular, MICE, MI, and DIMV were generally the most effective methods across different missing data rates, with MICE excelling at lower rates and MI and DIMV showing strong performance at higher rates.
For $r=0.2$, MICE achieved the lowest MSE of 0.169, indicating better imputation accuracy among the evaluated methods. MI followed with an MSE of 0.199, while Xgboost had the highest MSE at 0.209. In terms of MSE Shapley values, MICE again performed best with an MSE Shap of 0.015, slightly better than MI at 0.020. At $r=0.4$, DIMV demonstrated the best performance with an MSE of 0.280, outperforming MI (0.284) and SOFT-IMPUTE (0.285). The MSE Shapley values revealed that DIMV and SOFT-IMPUTE both achieved the lowest values of 0.026. For $r=0.6$, both MI and DIMV achieved the lowest MSE and MSE Shap of 0.399 and 0.038, respectively, highlighting their effectiveness in handling higher rates of missing data. After that, MI still outperformed other methods with MSE and MSE Shap of 0.509 and 0.047, respectively, as missing rates increased to the highest.
Interestingly, from both Table 2 and Table 3, one can see that a lower MSE may not imply a lower MSE Shap and vice versa. For example, in Table 3, at a missing rate of 0.4, the MSE of MICE is 0.307, which is higher than the MSE of missForest (0.298). However, the MSE Shap of MICE is 0.029, which is lower than of missForest (0.030).
(a) $r=0.2$
(b) $r=0.4$
(c) $r=0.6$
(d) $r=0.8$
Figure 13: MSE versus Shapley MSE plots on the diabetes dataset
6 Discussion
In summary, from the present results and the deep analysis in the previous section, we can observe several interesting key insights.
The global feature importance plots and the beeswarm plots clearly show that different imputation methods lead to varying Shapley value distributions. This shows that the choice of the imputation method can significantly affect the interpretability of the model.
Across different missing rates, XGBoost, without imputation, seems to have the most significant changes in the Shapley values. While Xgboost can train and predict directly on data with missing values, the MSE between Shapley values of XgBoost and the Original is the highest in most of the experiments, and even higher implementing XgBoost Regression after filling in missing values by mean imputation.
As the missing rate increases from 20% to 80% percent, the differences between the imputation methods become more pronounced. This indicates that the choice of the imputation method becomes increasingly critical as the number of missing data increases.
The results for the California and diabetes datasets show some differences, suggesting that the impact of imputation methods may be data set dependent. This highlights the importance of considering the characteristics of the data set when choosing an imputation method.
The MSE plots show that methods with lower imputation MSE do not necessarily preserve Shapley values better (as measured by Shapley MSE). This suggests a potential trade-off between the accuracy of the imputation and the maintenance of the original importance structure of the feature.
Mean imputation tends to significantly alter the importance of features, especially at higher missing rates. MICE and DIMV often show similar patterns, possibly due to the fact that MICE is based on regression and DIMV is based on a conditional Gaussian formula. MissForest and SOFT-IMPUTE sometimes preserve feature rankings better than simpler methods, but this is not consistent across all scenarios.
The variability in results between methods and missing rates underscores the need to evaluate imputation effects when using Shapley values for model interpretation.
The following discussion is structured around our result and the specific pitfalls that may arise due to incomplete understanding of the relationship between missing data, imputation methods, and Shapley values. We highlight how different approaches can lead to vastly different interpretations, how dataset characteristics and missing rates affect results, and the importance of considering both imputation accuracy and interpretability preservation.
Pitfall 1: Assume the neutrality of the imputation method. Our study reveals that different imputation methods can significantly alter Shapley values and, consequently, the interpretability of the model. For instance, mean imputation tends to distort feature importances, especially at higher missing rates, while methods like MICE and DIMV often show similar patterns. This underscores the importance of carefully considering the imputation method when using Shapley values for model explanation, as the choice can lead to vastly different interpretations of feature importance.
Pitfall 2: Overlooking data set dependency. We observed that the effects of imputation methods on Shapley values vary between data sets. For example, the California and diabetes datasets showed different patterns of importance of characteristics in different imputation methods. This highlights that dataset characteristics play a crucial role in determining the best imputation approach, and warns against applying a one-size-fits-all solution across different datasets.
Pitfall 3: Ignoring the impact of missing rate. Our results demonstrate that the impact of the choice of the imputation method becomes more pronounced as the missing rate increases. At lower missing rates, differences between methods are less stark, but as missing data increases, the choice of the imputation method becomes increasingly critical. This emphasizes the need for more careful consideration of imputation techniques when dealing with data sets with substantial missing data. In addition, a method may perform better than another one at low missing rates but perform worse than another one at high missing rates.
Pitfall 4: Focussing solely on imputation accuracy. We found a potential trade-off between the precision of the imputation and the preservation of the Shapley value. Methods that provide more accurate imputations do not necessarily better preserve the original Shapley values. This highlights a potential conflict between optimizing for prediction accuracy and maintaining interpretability and suggests that practitioners should consider both aspects when selecting an imputation method.
These pitfalls underscore the complex interaction between missing data handling, imputation methods, and model interpretability using Shapley values. Our findings highlight that there is no universal solution for handling missing data while preserving the interpretability of the model. Instead, the choice of method should be context-dependent, considering factors such as data set characteristics, missing data rates, and the specific requirements of the analysis. Moreover, our results emphasize the need for a holistic approach that balances imputation accuracy with the preservation of feature-importance structures. As machine learning models continue to be applied in critical domains, understanding and addressing these pitfalls becomes important to ensuring reliable and interpretable results.
7 Conclusion
Our comparative analysis revealed that the chosen imputation method should align with the specific characteristics of the data set and the objectives of the analysis. Practitioners should carefully consider the trade-offs between computational efficiency and the potential for bias introduction when selecting an imputation method. The study underscores the necessity of evaluating the effects of imputation as a critical step in the preprocessing pipeline, especially when Shapley values are used for model interpretation.
Future work should focus on extending this analysis to a broader range of datasets and machine learning models to further validate our findings. In addition, the development of new imputation methods is tailored for specific types of data, model structures, and explainability, especially in relation to Shapley value interpretations. Also, more research needs to be done on the direction of handling missing data directly, as this helps to avoid noises and bias introduced to the model by imputation. By addressing these challenges, we can improve the reliability of model interpretations and support more informed decision-making in the application of machine learning.
Appendix A Global feature importance plots on the diabetes dataset
(a) Original
(b) Xgboost
@@ Appendix A Global feature importance plots on the diabetes dataset
(a) Original
(b) Xgboost
@@
(a) Original
(b) Xgboost
@@
(a) Original
(b) Xgboost

### v2

@@ 5.2 Global feature importance plot analysis
To begin with, regardless of the missing rates, the global feature importance consistently highlights three main predictors of data: Latitude, Longitude, and Median Income (MedInc). However, their orders seem to change compared to the plot on the original data (LRO). For instance, at a missing rate of 0.2, LRO highlights Latitude the most, then Longitude and MedInc, while all the remaining methods highlight MedInc the most.
@@
Interestingly, across all missing rates, when comparing the global feature importance of three features (MedInc, Latitude, and Longitude) of LRO to the remaining methods, then XGBoost produces the most different plots. Specifically, XGBoost exhibits high global feature importance for those three features, with a range from $0.28$ to $0.69$. Meanwhile, for the six imputation methods, those values show a distribution similar to the results on the original data but at a lower range, from $0.05$ to $0.15$.
Additionally, as the missing rate increases from 0.2 to 0.8, there is a decrease in the global feature importance of the three mentioned features for the six comparison imputation methods, while XGBoost shows an upward trend. For example, for the MedInc feature, its global feature importance with the six imputation methods declines from around $0.15$ to $0.13$, whereas with XGBoost, it increases from $0.33$ to $0.69$.
(a) Original (LRO)
@@
(h) GAIN
5.3 Beeswarm plot analysis
While the global plot highlights overall feature importance, it doesn’t reveal how individual feature values impact predictions. The beeswarm plot fills this gap by showing the distribution and direction of Shapley values for each feature across all instances, giving insight into how specific feature values drive predictions up or down. Therefore, in this section, we deeply analyze the effects of feature value, missing rate, and handling missing strategy on the Shapley values by using the beeswarm plot. Note that the gray points in the beeswarm plots represent missing values.
5.3.1 Beeswarm plot for the California dataset
The beeswarm plot for the California datasets at missing rates $0.2,0.4,0.6,$ and $0.8$ are presented in Figures 5, 6, 7, and 8, respectively. In these figures, we focus on three key features (Latitude, Longitude, and MedInc) because the others show a relatively small impact.
Firstly, we examine the effect of feature values. Notably, feature values consistently influence the Shapley values across all comparison methods and missing rates. For example in Figure 5, at a missing rate r = 0.2, for the MedInc feature, its low value has a negative impact whereas its high value has a positive impact on model output regardless of all comparison methods.
Secondly, considering the results at each missing rate, we observed similar distributions in the Shapley values for the three key features on both LRO and six imputation methods. However, XGBoost shows a different distribution that is more skewed from zero. For instance, at a missing rate of 0.2 in Figure 5, the Shapley values in LRO for Latitude and Longitude range from $(-0.75)$ to $0.75$. Likewise, that range for six imputation methods is from around $(-0.5)$ to $0.5$. Meanwhile, XGBoost displays a different distribution from $(-1.2)$ to $0.6$ for Latitude and Longitude.
Finally, we focus on XGBoost without imputation. XGBoost seems to be stable across all missing rates. Especially, missing values always strongly affect the model output. For instance, at a missing rate of 0.4 in Figure 6, the missing values in the MedInc feature exhibit a strong positive impact, with Shapley values reaching as high as approximately $1.6$ or the missing values in the Longitude and Latitude features show a strong negative effect, with Shapley values dropping to around $(-1.2)$. In other words, for XGBoost without imputation, missing values tend to skew the distribution of Shapley values, either strongly negative or strongly positive.
(a) Original (LRO)
@@ 5.3.1 Beeswarm plot for the California dataset
(h) GAIN
5.3.2 Beeswarm plot for the Diabetes dataset
The beeswarm plot for the Diabetes dataset at missing rates $0.2,0.4,0.6,$ and $0.8$ are presented in Figures 9, 10, 11, and 12, respectively. A similar pattern in analyzing the impact of feature values on Shapley values can be seen, where feature values consistently influence Shapley values across all comparison methods and missing rates.
Regarding the distribution of Shapley values, across all missing rates, LRO and imputation methods exhibit a rather balanced distribution around the zero value. In contrast, the XGBoost model shows an increasingly skewed distribution as the missing rate rises. For example, focusing on the “bmi” feature in Figure 9, most of the Shapley values for LRO and imputation methods range from $(-0.25)$ to $0.25$, while for XGBoost, they range from $(-0.15)$ to $0.45$.
Considering XGBoost without imputation, as the missing rates increase, the Shapley values of missing values become increasingly skewed to one side. For instance, in Figure 12, at a missing rate of 0.8, the Shapley values for the features “bmi”, “bp”, “s5”, and “s6” are consistently positive, while those for “s2” and “s1” are predominantly negative.
(a) Original (LRO)
(b) XGBoost without imputation
@@ 5.3.2 Beeswarm plot for the Diabetes dataset
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
@@
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
@@
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
@@
(h) GAIN
5.4 MSE analysis
5.4.1 MSE analysis for regression tasks
The performances of XGBoost without imputation and various imputation methods are evaluated for regression on the California and Diabetes datasets at different missing rates, as shown in Table 2. Moreover, recall that the MSE represents the mean squared error between the true and predicted values in the test set labels. Also, MSE SHAP is the MSE between the Shapley values of the model fitted based on imputed data versus the Shapley values of the model fitted based on the original data.
Dataset $r$ Criteria XGBoost MI MICE DIMV missForest SOFT GAIN
California 0.2 MSE 0.223 0.254 0.239 0.239 0.174 0.263 0.317
MSE SHAP 0.354 0.175 0.077 0.054 0.061 0.096 0.217
0.4 MSE 0.238 0.250 0.271 0.268 0.239 0.262 0.300
MSE SHAP 0.345 0.129 0.082 0.067 0.093 0.100 0.142
0.6 MSE 0.240 0.253 0.270 0.270 0.243 0.263 0.308
MSE SHAP 0.362 0.134 0.090 0.076 0.097 0.110 0.149
0.8 MSE 0.231 0.243 0.263 0.260 0.233 0.254 0.293
MSE SHAP 0.346 0.136 0.090 0.073 0.098 0.106 0.151
Diabetes 0.2 MSE 0.193 0.188 0.161 0.150 0.170 0.155 0.216
MSE SHAP 0.063 0.019 0.016 0.016 0.017 0.016 0.021
0.4 MSE 0.309 0.298 0.332 0.275 0.320 0.304 0.358
MSE SHAP 0.125 0.029 0.028 0.027 0.030 0.028 0.033
0.6 MSE 0.359 0.366 0.386 0.371 0.445 0.349 0.427
MSE SHAP 0.142 0.038 0.039 0.035 0.042 0.037 0.044
0.8 MSE 0.549 0.535 0.563 0.574 0.660 0.555 0.603
MSE SHAP 0.224 0.049 0.051 0.050 0.058 0.048 0.055
Table 2: MSE on the test set labels and MSE of Shapley values for linear regression tasks on the California and Diabetes datasets.
Taking the results on the California dataset from Table 2, at $r=0.2$, missForest achieves the lowest MSE of 0.174, which is significantly lower than the second-best method, XGBoost, with an MSE of 0.223. However, when the missing rate exceeds 0.4, XGBoost consistently achieves the lowest MSE, followed closely by missForest. Interestingly, although missForest and XGBoost alternate in obtaining the lowest MSE, this does not correspond to the lowest MSE SHAP. Instead, DIMV demonstrates the best performance, indicating its robustness in preserving the explainability of the imputed data. For example, the MSE SHAP values for DIMV are 0.054, 0.067, 0.076, and 0.073, outperforming those of XGBoost, which are 0.354, 0.345, 0.362, and 0.346, as the missing rate increases from 0.2 to 0.8, respectively.
For the Diabetes dataset, overall, both MSE and MSE SHAP consistently increase as the missing rate increases. With a low missing rate (from 0.4 to lower), DIMV shows the best performance at both MSE and MSE SHAP. For example, at $r=0.2$, DIMV achieves the best MSE of 0.150, followed by SOFT with an MSE of 0.155. Also, DIMV, SOFT, and MICE demonstrate the best MSE SHAP, all of 0.016. As the missing rate increases, there is no consistency in the method that delivers the best results; instead, different methods alternately show their superiority. For instance, at $r=0.6$, SOFT outperforms the others in MSE with a value of 0.349, while DIMV achieves the best MSE SHAP at 0.035. At $r=0.8$, MI achieves the lowest MSE (0.535), while SOFT records the lowest MSE SHAP (0.048).
We, therefore, observe that different imputation methods excel in terms of prediction performance and explainability, underscoring the importance of selecting appropriate imputation methods based on specific criteria in the context of missing data analysis. The reason for the prominent performance of DIMV in terms of MSE SHAP could be that DIMV might better preserve the relationships between features and their contributions to the model’s predictions compared to the other approaches. In fact, DIMV is based on the Gaussian conditional formula with $L_{2}$ norm regularization. Meanwhile, missForest is a tree-based nonlinear approach, and SOFT-IMPUTE is a matrix decomposition approach to imputation. Also, the simplicity of DIMV and its regularization terms could reduce biases introduced by other imputation methods, leading to more accurate Shapley values.
5.4.2 MSE analysis for classification tasks
Dataset $r$ Criteria XGBoost MI MICE DIMV missForest SOFT GAIN
MNIST 0.2 MSE - 0.004 0.001 0.001 0.000 0.002 0.004
MSE SHAP 0.038 0.021 0.010 0.009 0.005 0.011 0.019
0.4 MSE - 0.009 0.004 0.003 0.002 0.005 0.008
MSE SHAP 0.074 0.034 0.020 0.018 0.011 0.023 0.034
0.6 MSE - 0.013 0.010 0.006 0.004 0.009 0.014
MSE SHAP 0.109 0.046 0.038 0.030 0.024 0.037 0.045
0.8 MSE - 0.017 0.017 0.012 0.011 0.027 0.020
MSE SHAP 0.142 0.053 0.053 0.044 0.041 0.053 0.052
Glass 0.2 MSE - 0.192 0.094 0.115 0.094 0.095 0.265
MSE SHAP 0.041 0.029 0.015 0.017 0.020 0.018 0.032
0.4 MSE - 0.782 1.617 0.624 0.612 0.578 1.319
MSE SHAP 0.186 0.080 0.063 0.057 0.071 0.062 0.0965
0.6 MSE - 0.706 1.258 0.675 0.707 0.623 0.924
MSE SHAP 0.149 0.064 0.059 0.050 0.060 0.050 0.070
Table 3: MSE on the test set inputs and MSE of Shapley values for classification tasks on the MNIST and Glass datasets.
For the classification tasks on the MNIST and Glass datasets, note that the MSE is calculated based on the test set inputs, and the results for the missing rate of 0.8 are unavailable because some rows have all features missing. Focusing on the MNIST dataset, missForest consistently delivers the best MSE and MSE SHAP across all missing rates. Specifically, the MSE values are 0 - 0.002 - 0.004 - 0.011, and the MSE SHAP values are 0.005 - 0.011 - 0.024 - 0.041 as the missing rate increases from 0.2 to 0.8. On the Glass dataset, at the missing rate of 0.2, MICE outperforms the others with an MSE of 0.094 and a MSE SHAP of 0.015. When the missing rate increases from 0.4 to 0.6, SOFT takes the lead with an MSE of 0.578 and 0.623, and MSE SHAP values of 0.062 and 0.050, respectively. One important observation is that the XGBoost classifier gives the highest MSE SHAP across all missing rates, thereby illustrating its lower ability to preserve the explainability of the imputed data compared to other methods.
6 Discussion
In summary, from the present results and the deep analysis in the previous section, we can observe several interesting key insights:
  • 1.
    The global feature importance plots and the beeswarm plots clearly show that different imputation methods lead to varying Shapley value distributions. This shows that the choice of the imputation method can significantly affect the explainability of the model.
  • 2.
    Across different missing rates, XGBoost, without imputation, seems to cause the most significant changes in the Shapley values. While XGBoost can train and predict directly on data with missing values, the MSE between Shapley values of XGBoost and the Original is the highest in most of the experiments.
  • 3.
    As the missing rate increases from 0.2 to 0.8, the differences between the imputation methods become more pronounced. This indicates that the choice of the imputation method becomes increasingly critical as the number of missing data increases.
  • 4.
    The results for the California and Diabetes datasets show some differences, suggesting that the impact of imputation methods may be data set dependent. This highlights the importance of considering the characteristics of the data set when choosing an imputation method. Such a research gap should be further investigated.
  • 5.
    The MSE plots show that methods with lower imputation MSE do not necessarily preserve Shapley values better (as measured by Shapley MSE). This suggests a potential trade-off between the accuracy of the imputation and the maintenance of the original importance structure of the feature. Interestingly, the results by [50] show that the methods that achieved the best MSE did not necessarily yield the most interpretable rule-based ML models, in the sense of model size and rule recovery.
  • 6.
    Mean imputation (MI) tends to significantly alter the importance of features, especially at higher missing rates.
  • 7.
    MICE and DIMV often show similar patterns, possibly due to the fact that MICE is based on regression and DIMV is based on a conditional Gaussian formula. MissForest and SOFT-IMPUTE sometimes preserve feature rankings better than simpler methods, but this is not consistent across all scenarios. GAIN does not show its superiority in all comparison datasets.
The variability in results between methods and missing rates underscores the need to evaluate imputation effects when using Shapley values for model interpretation for datasets with missing values. The following discussion is structured around our result and the specific pitfalls that may arise due to an incomplete understanding of the relationship between missing data, imputation methods, and Shapley values. We highlight how different approaches can lead to vastly different interpretations, how dataset characteristics and missing rates affect results, and the importance of considering both imputation accuracy and interpretability preservation:
  • 1.
    Pitfall 1: Assume the neutrality of the imputation method. Our study reveals that different imputation methods can significantly alter Shapley values and, consequently, the interpretability of the model. For instance, mean imputation and GAIN tend to distort feature importance, especially at higher missing rates, while methods like MICE and DIMV often show similar patterns. This underscores the importance of carefully considering the imputation method when using Shapley values for model explanation, as the choice can lead to vastly different interpretations of feature importance.
  • 2.
    Pitfall 2: Overlooking data set dependency. We observed that the effects of imputation methods on Shapley values vary between data sets. For example, the California and Diabetes datasets showed different patterns of importance of features in different imputation methods. This highlights that dataset characteristics play a crucial role in determining the best imputation approach, and warns against applying a one-size-fits-all solution across different datasets. Future work should address the question of what dataset characteristics are the most influential for the imputation.
  • 3.
    Pitfall 3: Ignoring the impact of missing rate. Our results demonstrate that the impact of the choice of the imputation method becomes more pronounced as the missing rate increases. At lower missing rates, differences between methods are less stark, but as missing data increases, the choice of the imputation method becomes increasingly critical. This emphasizes the need for more careful consideration of imputation techniques when dealing with data sets with substantial missing data. In addition, a method may perform better than another one at low missing rates but perform worse than another one at high missing rates, such as DIMV in the Diabetes dataset or MICE in the Glass dataset.
  • 4.
    Pitfall 4: Focusing solely on imputation accuracy. Our findings suggest a potential trade-off between imputation precision and the preservation of Shapley values. Methods that provide more accurate imputations do not necessarily better preserve the original Shapley values. This highlights a potential tension between optimizing for prediction accuracy and maintaining interpretability and suggests that practitioners should consider both aspects when selecting an imputation method.
These pitfalls underscore the complex interaction between missing data handling, imputation methods, and model interpretability using Shapley values. Our findings highlight that there is no universal solution for handling missing data while preserving the interpretability of the model. Instead, the choice of method should be context-dependent, considering factors such as data set characteristics, missing data rates, and the specific requirements of the analysis. Moreover, our results emphasize the need for a holistic approach that balances imputation accuracy with the preservation of feature-importance structures. As machine learning models continue to be applied in critical domains, understanding and addressing these pitfalls becomes important to ensuring reliable and interpretable results. As an additional suggestion for future work, it would be useful to develop explanation methods that are aware of whether some features in the training set have been imputed at training time and how often. Features that have been imputed more often are expected to be less reliable and, therefore, have relatively low importance.
7 Conclusion
Our comparative analysis revealed that the chosen imputation method should align with the specific characteristics of the data set and the objectives of the analysis. Practitioners should carefully consider the trade-offs between computational efficiency and the potential for bias introduction when selecting an imputation method. Moreover, our results revealed that imputation techniques that achieve the best accuracy do not necessarily produce the most accurate feature importance values. The study underscores the necessity of evaluating the effects of imputation as a critical step in the preprocessing pipeline, especially when Shapley values are used for model interpretation.
Future work should focus on extending this analysis to a broader range of datasets and machine learning models to validate our findings further. In addition, the development of new imputation methods is tailored for specific types of data, model structures, and explainability, especially in relation to Shapley value interpretations. Also, more research needs to be done on the direction of handling missing data directly, as this helps to avoid noises and bias introduced to the model by imputation. By addressing these challenges, we can improve the reliability of model interpretations and support more informed decision-making in the application of machine learning.
Appendix A Global feature importance plots on the Diabetes dataset
(a) Original (LRO)
(b) XGBoost without imputation
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(h) GAIN
(a) Original (LRO)
(b) XGBoost without imputation
(c) Mean Imputation
(d) MICE
(e) DIMV
(f) missForest
(g) SOFT-IMPUTE
(h) GAIN